OpenStudy (anonymous):

the density for manganese is 7.47g/cm3power what would be the mass of a 45cm3power piece of manganese

OpenStudy (anonymous):

\[ \frac{ 7.47g }{ 1 cm ^{3} } \times \frac{ 45cm ^{3} }{ 1 } = Answer\] Your answer is in grams